artlover03:

Solve *will post equation* , where is a real number. Simplify your answer as much as possible. If there is more than one solution, separate them with commas. If there is no solution, click on "No solution".

artlover03:

@snowflake0531 @jhonyy9

Florisalreadytaken:

Where is the equation?

artlover03:

Toast... sorry: \[w^2=36\]

jhonyy9:

\[w^{2} = 36 \] \[w_{1,2} = \sqrt{36}\]

jhonyy9:

hope you know 36 = 6*6 = 6^2 any idea now ?

artlover03:

exponets? or sqauring..?

jhonyy9:

\[w^{2} = 36 = 6^{2}\] \[w_{1,2}= \sqrt{6^{2}}= +/- ?\]

jhonyy9:

or an other way \[\sqrt{6^{2}}= +/- \left( 6^{2} \right)^\frac{ 1 }{ 2 } = ?\]

artlover03:

6,-6?

jhonyy9:

perfect

jhonyy9:

questions ?

artlover03:

no sir thank you :)

jhonyy9:

np anytime yw